# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v1

@@ 6 State Perception
Compared to passive perception, active perception adjusts the perspective to the areas of interest (Kroemer et al. (2021)). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning. The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (Han et al. (2022); Khan et al. (2022); Zhai et al. (2022)) and the scaling of ViTs has driven breakthrough capabilities for vision models (Dehghani et al. (2023)).
@@
Pre-trained visual representations have high transfer ability to policy learning (Xiao et al. (2022b); Yang et al. (2023b)), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (Karamcheti et al. (2023)). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (Majumdar et al. (2023)) and EmbCLIP (Khandelwal et al. (2022)) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough. At the same time, it is important to consider the potential of the diffusion model (Ze et al. (2023)).
The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. In the object point cloud modality, ULIP (Xue et al. (2023a)) and ULIP2 (Xue et al. (2023b)) employ contrastive learning to align features between point clouds and text-images. Point-BERT (Yu et al. (2022)) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. In the haptic modality, MOSAIC (Tatiya et al. (2023)) utilizes contrastive learning to train the haptic encoder.
@@
As for scene reconstruction, due to the foundation model’s robust open-set detection capability for objects, DFF (Shen et al. (2023)), CLIP-Fields (Shafiullah et al. (2022)) and LERF (Kerr et al. (2023)) employ CLIP to extract features from multi-view 2D images for NeRF (Mildenhall et al. (2021)) reconstruction. These features are then integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. 3D-LLM (Hong et al. (2023)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023)), endowing 3D features with semantic information. CLIP-NeRF (Wang et al. (2022)) integrates semantic features extracted by CLIP into NeRF reconstruction to change object textures during rendering.
Following Figure 6, representation learning facilitates the learning of the multimodal encoder, aligning features from various modalities to the text domain, aiding the model’s cognition, akin to how humans learn and think through words. The learning of the encoder promotes open-set perception tasks, such as open-set detection, open-set segmentation, and open-set scene reconstruction. When applied to specific scenarios, models often require fine-tuning. CLIP-Adapter (Gao et al. (2024)) offers a fine-tuning method. However, fine-tuning involves different techniques for different model architectures.
Figure 6: State Perception. The representation learning methods in state perception mainly include contrastive learning (Radford et al. (2021)), distillation-based learning (Caron et al. (2021)), and masked autoencoder learning (Radosavovic et al. (2023)). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). Multimodal representation learning focuses primarily on multimodal alignment (Xue et al. (2023b); Tatiya et al. (2023)). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation, and open-set scene reconstruction. For instance, SAM (Kirillov et al. (2023)) utilizes the MAE (He et al. (2022)), ViLD (Gu et al. (2021)) employs the CLIP (Radford et al. (2021)), and LERF (Kerr et al. (2023)) utilizes the CLIP (Radford et al. (2021)).
@@ 7 Policy
The policy model for outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next movement. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Code (VLAC), Vision-Language-Action-Target-Pose (VLATP), and Vision-Language-Action-Delta-Pose (VLADP) based on different output policies.
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, for a agent capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capabilities, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. In PROGPROMPT (Singh et al. (2023)), assertions are added to the generated code. When executing assertions, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task.
@@ 7.2 VLATP
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) and BC-Z (Jang et al. (2022)) demonstrate the capability of imitation learning in language-conditioned general manipulation. Due to the development of language foundation models, LEO (Huang et al. (2023a)) expands upon the language foundation model by incorporating modalities such as images and 3D point clouds and fine-tune manipulation datasets using the LoRA method, showcasing the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023a)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs.
@@
Figure 7: Foundation Models for Policy. 1) Vision-Language-Action-Code (VLAC) integrates task instruction, pre-written APIs, and example inputs into LLM, generating corresponding execution code. Then, when executing code, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task (Singh et al. (2023)). 2) Vision-Language-Action-Target-Pose (VLATP) without foundation models fine-turning inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation, and both the manipulated object and object segmentation images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D target pose (Huang et al. (2023b)). 3) Vision-Language-Action-Target-Pose (VLATP) with foundation models fine-tuning inputs task instruction and multimodal perception into a pre-trained model such as LLM or VLM after fine-tuning to output target pose (Huang et al. (2023a)). 4) Vision-Language-Action-Delta-Pose (VLADP) without fine-turning. LLM outputs the description of the object to be manipulated based on task instruction. VLM then outputs the BBox of the corresponding object in RGB. Next, segmentation is performed using VFM. 3D Value Map is updated based on depth information, generating delta pose (Huang et al. (2023c)). 5) Vision-Language-Action-Delta-Pose (VLADP) outputs delta pose and predicted observation image after action execution by inputting task instruction and multimodal perception into a pre-trained model after fine-tuning. The difference between VLADP and VLATP with fine-tuning lies in generating delta poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose. Target pose often requires subsequent motion planning. Outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining the image after the next movement, predicting both the next action and the hindsight image can improve the performance (Bousmalis et al. (2023)). 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (Di Palo et al. (2023)), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (Ma et al. (2023b)), while VLM judges the success of execution to provide criteria for RL rewards (Di Palo et al. (2023)).
@@ 7.3 VLADP
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (Reed et al. (2022)) and RT-1 (Brohan et al. (2022)) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Though there’s no evidence indicating the superiority of either input method, intuitively, solely observing through images may compromise stability by omitting proprioception. Building upon Gato, RoboCat (Bousmalis et al. (2023)) demonstrates that a large sequence model can learn unseen tasks through few-shot learning and proposes a simple but effective self-improvement process and also demonstrates that predicting both the next action and the hindsight image upon executing that action can enhance the performance. Building upon RT-1, RoboAgent (Bharadhwaj et al. (2023)) enhances model generalization and stability through data augmentation and action-chunking. MOO (Stone et al. (2023)) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
Utilizing pre-trained VLMs (Zhang et al. (2024)) for fine-tuning to construct RFMs is considered efficient. GR-1 (Wu et al. (2023a)) is initially pre-trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates sample data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). Therefore, RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses. Human intervention is added, enabling robots to adjust their trajectories promptly based on human input. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023b)) emerges, effectively enabling a robot manipulation policy with VLMs.
@@
Paper Robotics Dataset Collection Method Size Input Modality Architecture Hardware Output Modality Frequency Benchmark BC-Z (Jang et al. (2022)) BC-Z dataset (Jang et al. (2022)) Human Demonstrations1 Human Videos Robot Videos - Image Language or Video ResNet18 encoder + FiLM layers + a two-layer MLP 18 NVIDIA V100 GPUs Target Pose 10HZ Own benchmark (100 tasks in total) MOO (Stone et al. (2023)) RT-1 dataset (Brohan et al. (2022)) Self Collection Human Demonstrations  [^1] M Image Language RT-1 + OWL-ViT - Delta Pose - Own benchmark (1472 real world valuations) RT-1 (Brohan et al. (2022)) RT-1 dataset (Brohan et al. (2022)) Teleoperation 35M Image Language Conditioned EfficientNet + TokenLearnr + Decoder-only transformer - Delta Pose 3HZ Own benchmark (over 700 instructions) RT-2 (Brohan et al. (2023)) RT-1 dataset (Brohan et al. (2022)) Teleoperation 3/5/12/55B Image Language PaLI-X (Chen et al. (2022)), PaLM-E (Driess et al. (2023)) Multi-TPU Cloud Service Delta Pose 1-3HZ (Language Table Lynch et al. (2023)) Own benchmark RT-X (Padalkar et al. (2023a)) 60 individual datasets across 22 embodiments Human Demonstrations - Image Language PaLI-X (Chen et al. (2022)), PaLM-E (Driess et al. (2023)) Multi-TPU Cloud Service Delta Pose 3-10Hz Own benchmark (3600 evaluation trails) RT-H (Belkhale et al. (2024)) Diverse+Kitchen (Belkhale et al. (2024)) Human Demonstrations - Image Language PaLI-X (Chen et al. (2022)) - Delta Pose - Own benchmark (Kitchen tasks) LEO (Huang et al. (2023a)) CLIPort (Shridhar et al. (2021)) Human Demonstrations 1.3/7/13B Image 3D Point Cloud Language Decoder-only transformer 4 NVIDIA A100 GPUs Target Pose 1-5Hz Own benchmark (3 tasks for manipulation) RoboFlamingo (Li et al. (2023b)) CALVIN Bench(Mees et al. (2022)) Teleoperation 3/4/9B Image Language OpenFlamingo(Awadalla et al. (2023)) 8 NVIDIA Tesla A100 GPUs Delta Pose - CALVIN Bench(Mees et al. (2022)) Robocat (Bousmalis et al. (2023)) Meta-World (Yu et al. (2020)) DeepMind Control Suite (Tassa et al. (2018)) Self Collection Simulation Skill Library 364M/1.18B Image Language Proprioception Decoder-only transformer - Delta Pose Hindsight Image 10-20HZ RGB-Stacking Benchmark (Lee et al. (2021)) Roboagent (Bharadhwaj et al. (2023)) RoboSet(MT-ACT) (Bharadhwaj et al. (2023)) Teleoperation Skill Library - Image Language Proprioception Transformer One 2080Ti GPU Delta Pose 5Hz RL BENCH (James et al. (2020)) VIMA (Jiang et al. (2023)) VIMA dataset (Jiang et al. (2023)) Simulation 200M Image Language Decoder-only transformer 8 NVIDIA V100 GPUs Target pose - VIMA-BENCH (Jiang et al. (2023)) MIDAS (Li et al. (2023a)) VIMA dataset (Jiang et al. (2023)) Simulation 92M Image Language Decoder-only transformer 8 NVIDIA A100 GPUs Target pose - VIMA-BENCH (Jiang et al. (2023)) PaLM-E (Driess et al. (2023)) PaLM-E dataset (Driess et al. (2023)) Skill Library 12/22/562B Image Neural 3D Representations States of robot or objects Decoder-only transformer Multi-TPU Cloud Service Delta Pose Text 1-5Hz Language Table(Lynch et al. (2023)) Saycan(Ahn et al. (2022)) Gato (Reed et al. (2022)) Meta-World (Yu et al. (2020)) and other 15 datasets Simulation 1.2B Image Language Proprioception Decoder-only transformer TPU v3 Delta Pose 20HZ Meta-World (Yu et al. (2020)) GR-1 (Wu et al. (2023a)) Ego4D dataset (Grauman et al. (2022)) CALVIN Bench(Mees et al. (2022)) Self Collection Human Videos Teleoperation 195M Image Language Proprioception Decoder-only transformer - Delta Pose Image - CALVIN Bench(Mees et al. (2022)) QT-Opt (Kalashnikov et al. (2018)) QT-Opt dataset (Kalashnikov et al. (2018)) Simulation - Image Gripper Status Height Reinforcement Learning 10 NVIDIA P100 GPUs Target Pose - Own benchmark (28 test objects) RLS (Herzog et al. (2023b)) RLS dataset (Herzog et al. (2023b)) Skill Library (Simulation + Real World) - Image Proprioception CNN TPUv3 Target Pose - Own benchmark (9 scenarios) Q-Transformer (Chebotar et al. (2023b)) Q-Transformer dataset (Chebotar et al. (2023b)) Human Demonstrations Skill Library - Image Language Q-Transformer network - Target Pose 3HZ RT-1 (Brohan et al. (2022))

### v2

@@ 6 State Perception
Compared to passive perception, active perception adjusts the perspective to the areas of interest (Kroemer et al. (2021)). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning. The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (Han et al. (2022); Khan et al. (2022); Zhai et al. (2022)) and the scaling of ViTs has driven breakthrough capability for vision models (Dehghani et al. (2023)).
@@
Pre-trained visual representations have high transfer ability to policy learning (Xiao et al. (2022b); Yang et al. (2023c)), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (Karamcheti et al. (2023)). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (Majumdar et al. (2023)) and EmbCLIP (Khandelwal et al. (2022)) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough. At the same time, it is important to consider the potential of the diffusion model (Ze et al. (2023)). The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. For the object point cloud modality, ULIP (Xue et al. (2023a)) and ULIP2 (Xue et al. (2023b)) employ contrastive learning to align features between point clouds and text-images. Point-BERT (Yu et al. (2022)) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. In the haptic modality, MOSAIC (Tatiya et al. (2023)) utilizes contrastive learning to train the haptic encoder.
@@
As for scene reconstruction, due to the foundation model’s robust open-set detection capability for objects, DFF (Shen et al. (2023)), CLIP-Fields (Shafiullah et al. (2022)) and LERF (Kerr et al. (2023)) employ CLIP to extract features from multi-view 2D images for NeRF (Mildenhall et al. (2021)) reconstruction. These features are then integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. 3D-LLM (Hong et al. (2023)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023)), endowing 3D features with semantic information. CLIP-NeRF (Wang et al. (2022a)) integrates semantic features extracted by CLIP into NeRF reconstruction to change object textures during rendering.
Following Fig. 6, representation learning facilitates the learning of the multimodal encoder, aligning features from various modalities to the text domain, aiding the model’s cognition, akin to how humans learn and think through words. The learning of the encoder promotes open-set perception tasks, such as open-set detection, open-set segmentation, and open-set scene reconstruction. When applied to specific scenarios, models often require fine-tuning. CLIP-Adapter (Gao et al. (2024)) offers a fine-tuning method. However, fine-tuning involves different techniques for different model architectures.
Figure 6: Foundation Models for State Perception. The representation learning methods in state perception mainly include contrastive learning (Radford et al. (2021)), distillation-based learning (Caron et al. (2021)), and masked autoencoder learning (Radosavovic et al. (2023)). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). Multimodal representation learning focuses primarily on multimodal alignment (Xue et al. (2023b); Tatiya et al. (2023)). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation, and open-set scene reconstruction. For instance, SAM (Kirillov et al. (2023)) utilizes the MAE (He et al. (2022)), ViLD (Gu et al. (2021)) employs the CLIP (Radford et al. (2021)), and LERF (Kerr et al. (2023)) utilizes the CLIP (Radford et al. (2021)).
@@ 7 Policy
The policy model for outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next movement. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Code (VLAC), Vision-Language-Action-Target-Pose (VLATP), and Vision-Language-Action-Delta-Pose (VLADP) based on different output action.
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. In PROGPROMPT (Singh et al. (2023)), assertions are added to the generated code. When executing assertions, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task.
@@ 7.2 VLATP
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) and BC-Z (Jang et al. (2022)) demonstrate the capability of imitation learning in language-conditioned general manipulation. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023b)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. Xu et al. (2024) considers the task goal, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the target pose. However, the target pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point.
@@
Figure 7: Foundation Models for Policy. VLAC integrates task instruction, pre-written APIs, and example inputs into LLM. It generates corresponding execution code. When the code is executed, feedback from the environment is obtained. This feedback checks if the environment meets the task’s pre- and post-conditions (Singh et al. (2023)). 2) VLATP without foundation models fine-turning inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation, and both the manipulated object and object segmentation images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D target pose (Huang et al. (2023b)). 3) VLATP with foundation models fine-tuning inputs task instruction and multimodal perception into a pre-trained model such as LLM or VLM after fine-tuning to output target pose (Huang et al. (2023a)). 4) VLADP without fine-turning. LLM outputs the description of the object to be manipulated based on task instruction. VLM then outputs the BBox of the corresponding object in RGB. Next, segmentation is performed using VFM. 3D Value Map is updated based on depth information, generating delta pose (Huang et al. (2023c)). 5) VLADP outputs delta pose and the predicted observation image after action execution by inputting task instruction and multimodal perception into a pre-trained model after fine-tuning. The difference between VLADP and VLATP with fine-tuning lies in generating delta poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose. Target pose often requires subsequent motion planning. Outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining the image after the next movement, predicting both the next action and the hindsight image can improve the performance (Bousmalis et al. (2023)). 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (Di Palo et al. (2023)), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (Ma et al. (2023b)), while VLM can utilize prior knowledge to provide predicted action and sparse/dense reward, enhancing the effective exploration in reinforcement learning.(Ye et al. (2023a)).
@@ 7.3 VLADP
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (Reed et al. (2022)) and RT-1 (Brohan et al. (2022)) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Though there’s no evidence indicating the superiority of either input method, intuitively, solely observing through images may compromise stability by omitting proprioception. Building upon Gato, RoboCat (Bousmalis et al. (2023)) demonstrates that a large sequence model can learn unseen tasks through few-shot learning. It proposes a simple but effective self-improvement process. Additionally, it shows that predicting both the next action and the hindsight image after executing that action can enhance performance. Building upon RT-1, RoboAgent (Bharadhwaj et al. (2023)) enhances model generalization and stability through data augmentation and action-chunking. MOO (Stone et al. (2023)) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
Utilizing pre-trained VLMs (Zhang et al. (2024a)) for fine-tuning to construct RFMs is considered efficient. GR-1 (Wu et al. (2023a)) is initially pre-trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates sample data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). Therefore, RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses. Human intervention is added, enabling robots to adjust their trajectories promptly based on human input. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023c)) emerges, effectively enabling a robot manipulation policy with VLMs.
@@
Paper Robotics Dataset Collection Method Size Input Modality Architecture Hardware Output Modality Frequency Benchmark
BC-Z
(Jang et al. (2022))
BC-Z dataset
(Jang et al. (2022))
Human Demonstrations1
Human Videos
Robot Videos
Image
Language or Video
ResNet18 encoder +
FiLM layers +
a two-layer MLP
18 NVIDIA
V100 GPUs
Target Pose 10HZ
Own benchmark
(100 tasks in total)
MOO
(Stone et al. (2023))
RT-1 dataset
(Brohan et al. (2022))
+ Self Collection
Human Demonstrations  [^1] M
Image
Language
RT-1 + OWL-ViT - Delta Pose -
Own benchmark
(1472 real world valuations)
RT-1
(Brohan et al. (2022))
RT-1 dataset
(Brohan et al. (2022))
Teleoperation 35M
Image
Language
Conditioned EfficientNet +
TokenLearnr +
Decoder-only transformer
- Delta Pose 3HZ
Own benchmark
(over 700 instructions)
RT-2
((Brohan et al. (2023)))
RT-1 dataset
(Brohan et al. (2022))
Teleoperation 3/5/12/55B
Image
Language
PaLI-X
(Chen et al. (2022)),
PaLM-E
(Driess et al. (2023))
Multi-TPU
Cloud Service
Delta Pose 1-3HZ
Language Table
Lynch et al. (2023)
Own benchmark
RT-X
(Padalkar et al. (2023a))
60 individual datasets
across 22 embodiments
Human Demonstrations -
Image
Language
PaLI-X
(Chen et al. (2022)),
PaLM-E
(Driess et al. (2023))
Multi-TPU
Cloud Service
Delta Pose 3-10Hz
Own benchmark
(3600 evaluation trails)
RT-H
(Belkhale et al. (2024))
Diverse+Kitchen
(Belkhale et al. (2024))
Human Demonstrations -
Image
Language
PaLI-X
(Chen et al. (2022))
- Delta Pose -
Own benchmark
(Kitchen tasks)
LEO
(Huang et al. (2023a))
CLIPort
(Shridhar et al. (2021))
Human Demonstrations 1.3/7/13B
Image
3D Point Cloud
Language
Decoder-only transformer
4 NVIDIA
A100 GPUs
Target Pose 1Hz
Own benchmark
(3 tasks for manipulation)
RoboFlamingo
(Li et al. (2023c))
CALVIN Bench
(Mees et al. (2022))
Teleoperation 3/4/9B
Image
Language
OpenFlamingo
(Awadalla et al. (2023))
8 NVIDIA
Tesla A100 GPUs
Delta Pose -
CALVIN Bench
(Mees et al. (2022))
Robocat
(Bousmalis et al. (2023))
Meta-World
(Yu et al. (2020))
DeepMind Control Suite
(Tassa et al. (2018))
+Self Collection
Simulation
Skill Library
364M/1.18B
Image
Language
Proprioception
Decoder-only transformer -
Delta Pose
Hindsight Image
10-20HZ
RGB-Stacking Benchmark
(Lee et al. (2021))
Roboagent
(Bharadhwaj et al. (2023))
RoboSet(MT-ACT)
(Bahl et al. (2023))
Teleoperation
Skill Library
Image
Language
Proprioception
Transformer One 2080Ti GPU Delta Pose 5Hz
RL BENCH
(James et al. (2020))
VIMA
(Jiang et al. (2023))
VIMA dataset
(Jiang et al. (2023))
Simulation 200M
Image
Language
Decoder-only transformer
8 NVIDIA
V100 GPUs
Target pose -
VIMA-BENCH
(Jiang et al. (2023))
MIDAS
(Li et al. (2023b))
VIMA dataset
(Jiang et al. (2023))
Simulation 92M
Image
Language
Decoder-only transformer
8 NVIDIA
A100 GPUs
Target pose -
VIMA-BENCH
(Jiang et al. (2023))
PaLM-E
(Driess et al. (2023))
PaLM-E dataset
(Driess et al. (2023))
Skill Library 12/22/562B
Image
Neural 3D Representations
States of robot or objects
Decoder-only transformer
Multi-TPU
Cloud Service
Delta Pose
Text
1-5Hz
Language Table
(Lynch et al. (2023))
Saycan
(Ahn et al. (2022))
Gato
(Reed et al. (2022))
Meta-World
(Yu et al. (2020))
and other 15 datasets
Simulation 1.2B
Image
Language
Proprioception
Decoder-only transformer TPU v3 Delta Pose 20HZ
Meta-World
(Yu et al. (2020))
GR-1
(Wu et al. (2023a))
Ego4D dataset
(Grauman et al. (2022))
CALVIN Bench
(Mees et al. (2022))
+Self Collection
Human Videos
Teleoperation
195M
Image
Language
Proprioception
Decoder-only transformer -
Delta Pose
Image
CALVIN Bench
(Mees et al. (2022))
QT-Opt
(Kalashnikov et al. (2018))
QT-Opt dataset
(Kalashnikov et al. (2018))
Simulation -
Image
Gripper Status
Height
Reinforcement Learning
10 NVIDIA
P100 GPUs
Target Pose -
Own benchmark
(28 test objects)
RLS
(Herzog et al. (2023b))
RLS dataset
(Herzog et al. (2023b))
Skill Library
(Simulation + Real World)
Image
Proprioception
CNN TPUv3 Target Pose -
Own benchmark
(9 scenarios)
Q-Transformer
(Chebotar et al. (2023b))
Q-Transformer dataset
(Chebotar et al. (2023b))
Human Demonstrations
Skill Library
Image
Language
Q-Transformer network - Target Pose 3HZ
RT-1
(Brohan et al. (2022))